Banking fees have received much attention during the recent economic recession as banks look for ways to recover from the crisis. A sample of 32 customers paid an average fee of ​$12.11 per month on their​ interest-bearing checking accounts. Assume the population standard deviation is ​$1.65

Complete parts a and b below.

a. Construct a 90​% confidence interval to estimate the average fee for the population.The 90​% confidence interval has a lower limit of blank and an upper limit of.

​(Round to the nearest cent as​ needed.)
Answer #1

Solution :

Given that,

Point estimate = sample mean = \bar x = $12.11

Population standard deviation = \sigma = $1.65

Sample size = n = 32

At 90% confidence level the z is ,

  = 1 - 90% = 1 - 0.90\alpha = 0.10

\alpha / 2 = 0.10 / 2 = 0.05

Z\alpha/2 = Z0.05 = 1.645

Margin of error = E = Z\alpha/2* (\sigma /\sqrtn)

= 1.645 * (1.65 / \sqrt 32)

= 0.48

At 99% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

12.11 - 0.48 < \mu < 12.11 + 0.48

11.63 < \mu < 12.59

lower limit = 11.63

upper limi = 12.59
Answer #2

Banking fees have received much attention during the recent economic recession as banks look for ways to recover from the crisis. A sample of

customers paid an average fee of

$

per month on their interest-bearing checking accounts. Assume the population standard deviation is

$.

Complete parts a and b below


answered by: t